OpenStudy (anonymous):

Please help! I need to find the solution to the system x' = 6y, y' = -5x, with the initial conditions being x(0) = 1, y(0)= 5. I believe I have to do eigenvalues, but my eqns aren't coming out right.

OpenStudy (jamesj):

What eigenvalues have you found?

OpenStudy (anonymous):

I think I have found lambda equals +/- sqrt(30), based on my matrix: [0- lambda, 6; -5, 0-lambda]

OpenStudy (jamesj):

+- sqrt(30) i They are imaginary.

OpenStudy (anonymous):

OH. my goodness. Math error! Thanks!

OpenStudy (anonymous):

But---

OpenStudy (jamesj):

and hence the eigenvectors will also be complex. To deal with this, use the same trick you use when the roots of the characteristic equation are imaginary: take sums to find trig functions cos and sin

OpenStudy (anonymous):

I want to double check my method: okay, so, I find the values, then the vectors, and then I set up an eqn that looks some thing like x(t) = c1v1e^(eigenvector)t...?

OpenStudy (anonymous):

oh not, I replace the e^(eigenvector) with (cost +isint)

OpenStudy (jamesj):

call the eigenvalues l1 and l2, with corresponding eigenvectors v1 and v2. Then the general solution is c1.v1.e^(l1.t) + c2.v2.e^(l2.t) where c1 and c2 are constants.

OpenStudy (anonymous):

but what happens to the (cost + isint)?

OpenStudy (jamesj):

Yes, change basis and have two new eigenfunctions: 1/2 . ( e^(l1.t) + e^(l2.t) ) = cos(l1.t) 1/2i . ( e^(l1.t) - e^(-l2.t) ) = sin(l1.t) remember that l1 = -l2

OpenStudy (amistre64):

never tried these before but i did just watch a video by arthur mattuck tho. x' = 6y +0x y' = 0y - 5y would our matrix be: 6 0 0 -5 or am i going astray there?

OpenStudy (jamesj):

@amistre: your coefficients are in the wrong place

OpenStudy (jamesj):

x' = 0x + 6y etc.

OpenStudy (amistre64):

we have to start with the x part?

OpenStudy (jamesj):

The column vector of the functions x and y have to be in the same order on both sides of the equation

OpenStudy (anonymous):

yep. And so now I finally have my vector [5; -30i]

OpenStudy (anonymous):

nope, i think it should be [5, 30i] for complex conjugates

OpenStudy (anonymous):

nvm keep it the way is was

OpenStudy (anonymous):

this is wrong...

OpenStudy (anonymous):

help :/ I don't know what I am doing any more :C I know that my eigen values are +/- 30i, and when I try to find the eigenvectors, I find -30i*a + 5*b = 0 and -5*a - 30i*b = 0

OpenStudy (anonymous):

I don't thing the vector v = [5, -30i] is correct; and this affects my eqns for x and y

OpenStudy (amistre64):

\[\begin{vmatrix}6&0\\0&-5\end{vmatrix}\to \lambda^2+\lambda-30=0;\ \lambda={-6,5} \] \[\begin{vmatrix}0&6\\-5&0\end{vmatrix}\to \lambda^2+ 30=0;\ \lambda={\pm i\sqrt{30}} \] so i spose it does matter which column is where

OpenStudy (anonymous):

Yeah, the x values are far left and the y values are far right. So your bottom one is correct. Sorry, I didn't put sqrt() around my 30

OpenStudy (anonymous):

I am making my vectors again

OpenStudy (anonymous):

so it should be v = [-sqrt(30)i; 6]

OpenStudy (anonymous):

hello, there anyone else who can help?